OpenStudy (anonymous):

What are the zeroes of the function? What are their multiplicities? = x4 – 4x3 + 3x2 (1 point)

OpenStudy (anonymous):

f(x)= x4 – 4x3 + 3x2

OpenStudy (anonymous):

x^4-4x^3+3x^2=0 Solve for x to find zeros. CAN U?

OpenStudy (anonymous):

to find zeros f(x)=0 x^4-4x^3+3x^2=0 SOLVE FOR X.

OpenStudy (anonymous):

what is " ^" that symbol supposed to be?

OpenStudy (anonymous):

and I don't know how to solve for x

OpenStudy (anonymous):

try to factor f(x)

OpenStudy (phi):

^ means raise to the power saying x4 is ambiguous (it could mean x times 4). so say x^4 for \(x^4\)

OpenStudy (anonymous):

so would the answer be The numbers –1 and –3 are zeroes of multiplicity 2; the number 0 is a zero of multiplicity 1?

OpenStudy (anonymous):

or is that completely wrong?

OpenStudy (anonymous):

i seriously have no idea how to do this... :/

OpenStudy (phi):

do you know how to "factor out" x^2 from \[f(x)= x^4 – 4x^3 + 3x^2\] use this rule: (ab + ac) = a(b+c) (this is the opposite of distribute) you get \[ x^2(x^2-4x+3) \]

OpenStudy (phi):

multiplicity means how many times a root is repeated the roots are the numbers that make f(x) be 0 when you have \[ x^2(x^2-4x+3)=0\] which is the same as \[ x\cdot x\cdot (x^2-4x+3) =0\] if you replace x with 0, the whole thing becomes 0. so x=0 is a root because x is repeated (you have x times x or x^2), 0 has multiplicity 2 now you have to find the 2 roots that make \(x^2-4x+3=0\)

OpenStudy (anonymous):

would it be The numbers 0 and 1 are zeroes of multiplicity 2; the number 3 is a zero of multiplicity 1?

OpenStudy (anonymous):

The number 0 is a zero of multiplicity 2; the numbers 1 and 3 are zeroes of multiplicity 1

OpenStudy (phi):

no. First, this polynomial has a highest power of 4. that means you will have 4 roots as explained above, 0 is repeated, so 2 of the roots are 0,0 or just 0 with multiplicity 2 there are 2 more roots to find

OpenStudy (phi):

he number 0 is a zero of multiplicity 2; the numbers 1 and 3 are zeroes of multiplicity 1 Yes. I assume you guessed? you can check by replacing x with 1 in the equation. If you get 0 then you know 1 is a root. to solve for these roots, you factor x^2 -4x+3=0

OpenStudy (phi):

See http://www.khanacademy.org/math/algebra/polynomials/v/factoring-quadratic-expressions for how to do this if you don't know how